Mark is 10 years younger than Larry. Larry's age in 8 years will exceed twice Mark's age 3 years ago by 4 years. How old is each now.
Answer by oscargut(891)

Mark is 10 years younger than Larry. Larry's age in 8 years will exceed twice Mark's age 3 years ago by 4 years. How old is each now.
Let Mark's age = M and Larry's age = L
Mark is 10 years younger than Larry. then
1) M+10=L
L+8 = Larry's age in 8 years
2(M-3) = twice Mark's age 3 years ago
Larry's age in 8 years will exceed twice Mark's age 3 years ago by 4 years.
then L+8=2(M-3)+4 then L+8=2M-2 then
2) L=2M-10
Then using 1) and 2) M+10=2M-10 then -M=-20 then M=20
and using 1) L=30
R: Larry's age =30 and Mark's age =20